# Deep Optimal Transport for Domain Adaptation on SPD Manifolds

### v4

@@ VI DISCUSSIONS
We claim that Theorem 3 with cost function $c(x,\bar{x})=||x-\bar{x}||_{\ell^{2}}^{2}$ is equilvant to [6, Theorem 3.1]. It was first observed and proved by Yair et al.[10] who use the matrix vectorization operator to establish an isometry between the affine transformation and the Bi-Map transformation. We follow their arguments and revise a bit of their proof.
@@
[6, Theorem 3.1]  Let $\mu$ and $\nu$ be two discrete distributions on $\mathbb{R}^{n}$ with $N$ Diracs. Given a strictly positive definite matrix $A$, bias weight $b\in\mathbb{R}^{n}$, and source samples $\{x_{i}\}_{i=1}^{N}\sim\mu$, suppose target samples $\bar{x}_{i}:=Ax_{i}+b$, for $i=1,...,N$, and the weights in both source and target distributions’ empirical distributions are $1/N$. Then, transport $T(x_{i})=Ax_{i}+b$ is the solution to the OT Problem provided with a cost function $c(x,\bar{x}):=||x-\bar{x}||_{\ell^{2}}^{2}$.
To expose the relationship between the Bi-Map transformation in Theorem 3 and affine transformation in Theorem 4, we introduce the matrix vectorization operator $vec(\cdot)$ to stack the column vectors of matrix $A=(a_{1}\big{|}a_{2}\big{|}\cdots\big{|}a_{n})$ below on another as follows,
@@
According to Lemma 5 and 6, the Bi-Map transformation $W\cdot S\cdot W^{\top}$ turns to be affine transformation $A=W\otimes W$, and thus Theorem 3. Hence, Theorem 3 endowed with the cost function $c\big{(}vec(x),vec(y)\big{)}=||vec(x)-vec(y)||_{\ell^{2}}^{2}$ is equilvant to Theorem 4.
VI-B Parallel Transport
@@
Keep in mind that DOT seeks a neural networks-based transformation $\varphi:X_{S},X_{T}\longmapsto Z\subset\mathcal{S}_{++}$ such that $\mathbb{P}_{S}(\varphi(X_{S}))=\mathbb{P}_{T}(\varphi(X_{T}))$ and $\mathbb{Q}_{S}(Y_{S}|\varphi(X_{S}))=\mathbb{Q}_{T}(Y_{T}|\varphi(X_{T}))$, where $\varphi(X_{S})$ and $\varphi(X_{T})$ are transformation $\varphi$ on the source and target feature spaces. Neural networks-based transformation $\varphi$ extracts discriminative information from feature spaces. After the extraction, it is reasonable to view the feature space of the source domain as the same as it of the target domain, as they will be transformed under neural networks with the same weights, so are their transformed feature spaces, i.e., $X_{T}=X_{S}$ and $\varphi(X_{T})=\varphi(X_{S})\subset Z$, and thus we utilize the Log-Euclidean metric to formulate the space of covariance matrice.
VI-C Assumption 3
In this section, we will justify that Assumption 3 is reasonable from the numerical results. Table III records average Log-Euclidean distances in the same frequency band, and diagonal numbers are consistently the smallest among all the numbers in each row for the two datasets. This suggests that the mapping shifts the center of each frequency band in the source domain to the corresponding frequency band center in the target domain. It aligns with our intuition that the difference between the same frequency bands should be smaller since they capture similar signal characteristics in the same frequency range, and the difference between different frequency bands tends to be larger.
The procedure of calculating the average Log-Euclidean distance is specifically given as follows: assuming source and target samples are represented in a format of [trials, frequency bands, channels, channels], with frequency bands fixed (in the 2nd dimension), we first calculate the Fréchet mean by averaging all covariance matrices along the trial dimension. This results in the shape of [frequency bands, channels, channels]. Next, we directly compute the average Log-Euclidean distance for each pair of frequency bands.
@@ VI-C Assumption 3
TABLE III: Average Log-Euclidean distances between the Fréchet means of EEG spatial covariance matrices generated from different frequency bands in the source and target domains for the KU (54 subjects), BNCI2014001 (9 subjects), and BNCI2015001 (12 subjects) datasets. The shortest average Log-Euclidean distance in each row is highlighted in boldface.
@@
VII CONCLUSIONS
This study advances the OT-DA framework to SPD manifolds equipped with the Log-Euclidean metric and introduces a geometric deep learning approach to address real-world domain adaptation challenges with SPD matrix-valued data. Specifically, it handles cross-session variabilities in EEG-based motor imagery classification. Experiments on three publicly available EEG datasets validate the proposed approach’s effectiveness. The results highlight its practical potential and advantages in managing complex tasks involving domain adaptation in EEG data analysis.
@@ Acknowledgment
This study is supported under the RIE2020 Industry Alignment Fund–Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contributions from the industry partner(s). This study is also supported by the RIE2020 AME Programmatic Fund, Singapore (No. A20G8b0102).

### v5

@@ 6 DISCUSSIONS
We claim that Theorem 3 with cost function $c(x,\bar{x})=||x-\bar{x}||_{\ell^{2}}^{2}$ is equilvant to [5, Theorem 3.1]. It was first observed and proved by Yair et al.[11] who use the matrix vectorization operator to establish an isometry between the affine transformation and the BiMap transformation. We follow their arguments and revise a bit of their proof.
@@
[5, Theorem 3.1]  Let $\mu$ and $\nu$ be two discrete distributions on $\mathbb{R}^{n}$ with $N$ Diracs. Given a strictly positive definite matrix $A$, bias weight $b\in\mathbb{R}^{n}$, and source samples $\{x_{i}\}_{i=1}^{N}\sim\mu$, suppose target samples $\bar{x}_{i}:=Ax_{i}+b$, for $i=1,...,N$, and the weights in both source and target distributions’ empirical distributions are $1/N$. Then, transport $T(x_{i})=Ax_{i}+b$ is the solution to the OT Problem provided with a cost function $c(x,\bar{x}):=||x-\bar{x}||_{\ell^{2}}^{2}$.
To expose the relationship between the BiMap transformation in Theorem 3 and affine transformation in Theorem 4, we introduce the matrix vectorization operator $vec(\cdot)$ to stack the column vectors of matrix $A=(a_{1}\big{|}a_{2}\big{|}\cdots\big{|}a_{n})$ below on another as follows,
@@
According to Lemma 5 and 6, the BiMap transformation $W\cdot S\cdot W^{\top}$ turns to be affine transformation $A=W\otimes W$, and thus Theorem 3. Hence, Theorem 3 endowed with the cost function $c\big{(}vec(x),vec(y)\big{)}=||vec(x)-vec(y)||_{\ell^{2}}^{2}$ is equilvant to Theorem 4.
6.2 Parallel Transport Under Log-Euclidean Metric
@@
The DOT framework seeks a neural networks-based transformation $\varphi:X_{S},X_{T}\longmapsto Z\subset\mathcal{S}_{++}$ such that $\mathbb{P}_{S}(\varphi(X_{S}))=\mathbb{P}_{T}(\varphi(X_{T}))$ and $\mathbb{Q}_{S}(Y_{S}|\varphi(X_{S}))=\mathbb{Q}_{T}(Y_{T}|\varphi(X_{T}))$, where $\varphi(X_{S})$ and $\varphi(X_{T})$ are transformation $\varphi$ on the source and target feature spaces. Neural networks-based transformation $\varphi$ extracts discriminative information from feature spaces. After the extraction, it is reasonable to view the feature space of the source domain as the same as it of the target domain, as they will be transformed under neural networks with the same weights, so are their transformed feature spaces, i.e., $X_{T}=X_{S}$ and $\varphi(X_{T})=\varphi(X_{S})\subset Z$, and thus we utilize the Log-Euclidean metric to formulate the space of covariance matrice.
6.3 Pseudo-Label-Driven Conditional Distribution Adaptation
In this study, we adopt a pseudo-labeling strategy inspired by the JDA framework [12] to facilitate the alignment of conditional distributions between source and target domains. Specifically, we assign pseudo-labels to the unlabeled target domain samples using a base classifier trained on labeled source domain data. While discrepancies in both marginal and conditional distributions may introduce labeling noise, this strategy provides a practical foundation for conditional alignment in the absence of target labels. For EEG classification tasks, we employ Graph-CSPNet as the base classifier, as it outperforms the best in this setting [24]. Importantly, this pseudo-labeling mechanism not only enables conditional distribution alignment but also improves cross-domain adaptation by capturing domain-specific discriminative features.
6.4 Time Complexity in Optimal Transport
The computational complexity of the discrete OT problem poses significant challenges, especially for high-dimensional or large-scale datasets. Traditional methods for the discrete OT problem—such as the simplex algorithm or the Hungarian method—often require $O(N^{3})$ time when solving an $N\times N$ assignment problem, limiting their application to small-scale scenarios [46]. Entropy-regularized approaches, such as the Sinkhorn algorithm, reduce the computational burden by achieving a per-iteration complexity of $O(KNM)$ (which becomes $O(KN^{2})$ when $N=M$) over $K$ iterations [47]. For high-dimensional settings, techniques like sliced Wasserstein distances—which project data onto $L$ random one-dimensional subspaces—have a computational complexity of $O(LN\log N)$ [48].
In practical applications, trade-offs are essential. While exact methods are suitable for small, controlled problems, in EEG classification, the decision to use Wasserstein barycenters under the Log-Euclidean metric as intermediate reference distributions with neural networks is guided by specific application requirements. The Barycenter under the Log-Euclidean metric, which has been simplified as the arithmetic mean of their tangent representation, proves particularly beneficial in multi-source domain adaptation scenarios (for example, cross-session BCI calibration) by reducing computational overhead and enhancing stability.
6.5 Average Log-Euclidean Distance in Assumption 3
This section provides empirical validation for the reasonableness of Assumption 3 through numerical analysis. Table 3 presents the average Log-Euclidean distances between frequency bands across domains. Notably, the diagonal entries—representing distances between the same frequency bands in the source and target domains—consistently exhibit the smallest values within each row for both datasets. This observation validates that our mapping successfully aligns the center of each source domain frequency band with its corresponding target domain counterpart. The results align with the intuitive expectation that covariance structures within the same frequency band (e.g., alpha or beta bands) should show smaller divergences due to shared neurophysiological characteristics, whereas comparisons across different frequency bands naturally result in larger distances.
The computation protocol is as follows: Given source and target domain samples structured as [trials × frequency bands × channels × channels], we first compute the Fréchet mean for each frequency band by averaging the covariance matrices across trials, resulting in a [frequency bands × channels × channels] representation. We then compute pairwise Log-Euclidean distances between all frequency band pairs, where intra-band distances (on the diagonal) reflect the accuracy of alignment, while inter-band distances (off-diagonal) capture the divergence between different frequency bands.
@@ 6.5 Average Log-Euclidean Distance in Assumption 3
Table 3: Average Log-Euclidean distances between the Fréchet means of EEG spatial covariance matrices generated from different frequency bands in the source and target domains for the KU (54 subjects), BNCI2014001 (9 subjects), and BNCI2015001 (12 subjects) datasets. The columns and rows are the sequence of frequency bands. The shortest average Log-Euclidean distance in each row is highlighted in boldface.
KU
@@
7 CONCLUSIONS
This study extends the OT-DA framework to operate on SPD manifolds equipped with the Log-Euclidean metric. We propose a novel geometric deep learning approach for domain adaptation with SPD matrix-valued data by aligning both marginal and conditional distributions. The method specifically addresses variability between sessions in EEG-based motor imagery classification - a long-standing challenge in brain-computer interface research, where changes in signal distributions between sessions often degrade model performance. Experiments on three public EEG datasets (KU, BNCI2014001, BNCI2015001) validate the effectiveness of the proposed method, showing that it can effectively reduce distribution shifts between sessions.
Despite these contributions, the approach still faces several limitations. One significant challenge is the sensitivity to errors in pseudo-labels, even though the alignment of conditional distributions is stabilized using sufficient statistics. Early-stage label predictions, which often have around 70% accuracy in cross-session EEG settings, can hinder adaptation due to error propagation. This issue limits the overall effectiveness of the adaptation process. Furthermore, while transfer learning methods show measurable effectiveness in improving classification performance, domain change is not the only determinant of performance. As a result, the improvements from transfer techniques are often modest. A major limitation is the lack of a robust quantitative framework to measure the intensity of domain shift and its systematic relationship with classifier performance. This gap prevents a comprehensive evaluation of the true efficacy of the proposed methodology. Future work should focus on addressing these challenges, particularly by enhancing the robustness of pseudo-labeling through uncertainty-aware alignment strategies.
@@ Acknowledgment
This work was supported under the RIE2020 Industry Alignment Fund–Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contributions from industry partner(s); This work was also supported by the RIE2020 AME Programmatic Fund, Singapore (No. A20G8b0102).
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author(s) used ChatGPT 4 (OpenAI) in order to improve the English grammar of this manuscript. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.